OpenStudy (ezraweise):

What is the solution to the equation? A. z = –44 B. z = –20 C. z = 20 D. z = 44

OpenStudy (whpalmer4):

\[3-\frac{Z}{4} = 8\]Start by subtracting 3 from both sides, what do you get?

OpenStudy (ezraweise):

from 3 an 8?

OpenStudy (whpalmer4):

\[3-3 - \frac{Z}{4} = 8-3\]Can you simplify that?

OpenStudy (ezraweise):

ya

OpenStudy (ezraweise):

ANSWER ASAP PLEASE ANYONE PLEASE!!!!!!!!

OpenStudy (whpalmer4):

well, what do you get when you do?

OpenStudy (ezraweise):

anyone

OpenStudy (whpalmer4):

no one is going to give you the answer, but we'll help you work it out for yourself.

OpenStudy (ezraweise):

8-3

OpenStudy (whpalmer4):

no...what is \[3-3-\frac{Z}{4}\]after you simplify it?

OpenStudy (ezraweise):

0

OpenStudy (whpalmer4):

what happens to \[-\frac{Z}{4}\]?

OpenStudy (ezraweise):

hold be back in a sec

OpenStudy (ezraweise):

k back

OpenStudy (ezraweise):

\[Help\]

OpenStudy (ezraweise):

answer asap

OpenStudy (whpalmer4):

like I said, what happened to -Z/4 when you simplified? You just eliminated it, why?

OpenStudy (ezraweise):

@Eli_Moses come

OpenStudy (whpalmer4):

You misunderstand the purpose of this site. It is NOT to supply you with answers.

OpenStudy (eli_moses):

whazzap

OpenStudy (ezraweise):

answer this

OpenStudy (ezraweise):

yes some times it is @whpalmer4

OpenStudy (ezraweise):

and sometimes its not :-////

OpenStudy (whpalmer4):

Once again: \[3 - 3 - \frac{Z}{4}\]the 3's combine to 0. Don't you still have something left over?

OpenStudy (ezraweise):

leave report me all you want @whpalmer4 your blocked now there is nothing you can do SORRY :-////

OpenStudy (whpalmer4):

You must be confused about what blocking does.

OpenStudy (ezraweise):

oh LEAVE!